determine the equation with the line with x-intercept -1/3 and y-intercept -1
Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(75887)

determine the equation with the line with x-intercept -1/3 and y-intercept -1
---------------
You have two points:: (-1/3,0) and (0,-1)
-------
Form: y = mx + b
-------
m = (-1-0)/(0--(1/3)) = -1/(1/3) = -3
--------
Substitute for x,y, and m ; Solve for "b":
-1 = -3*0 + b
b = -1
----------
Equation:
y = -3x - 1
==================
Cheers,
Stan H.

Answer by lwsshak3(11628)

determine the equation with the line with x-intercept -1/3 and y-intercept -1
***
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
For given line:
y=mx-1
solve for m using coordinates of x-intercept,(-1/3,0)
0=m*-1/3-1
-m/3=1
m=-3
equation:
y=-3x-1